A chemistry student must write down in her lab not
Answer #1

Empty cyclinder weight = 1.5 g

After addition of sodium hydroxide into cylinder weight = 13.195 g

sodium hydroxide weight, which was put in cyclinder = 13.195-1.5 = 11.695 g

Total volume of the solution = 127.53 ml

Molarity = (Wt/MWt)x(1000/V in ml) = (11.695/40)x(1000/127.53) = 11695/5101.2 = 2.292 g/mL

The concentration of solution of sodium hydroxide = 2.292 g/mL